# Zero-order Control Barrier Functions for Sampled-Data Systems with State and Input Dependent Safety Constraints

### v1

@@ I Introduction
However, there exist limitations for the conventional CBF approach. Safety constraints must be independent of the system input, and the function defining the safe set must have relative degree $1$. Otherwise, the CBF condition is either not well-defined or becomes non-binding. In order to accommodate these cases, several extensions to conventional CBFs have been proposed. Notably, exponential CBFs [8] and high-order CBFs [9, 10] are proposed for addressing the high-relative degree issue. However, checking whether the recursively defined safety function is a valid CBF in this approach is difficult in general, and the performance is highly sensitive to the chosen coefficients. Alternatively, back-stepping CBFs [11] define barrier functions based on virtual inputs and tracking errors. However, back-stepping can encounter an “explosion of terms” problem when the relative degree is large. On the other hand, when the safety constraints depend both on system states and system inputs, integral CBFs [12] and control-dependent CBFs [13] extend the system dynamics via a new virtual input, treat the original state and input as the new state, and then apply the conventional CBF approach for this extended system. For all these extensions, the relation between the input and its effects on the safety constraints is not explicit.
Digital implementation of a controller designed for continuous-time systems usually involves a zero-order hold, resulting in a sampled-data system formulation: the state evolves in continuous-time, while the control input is updated at discrete sampling instants, and held constant between those instants. Many works have investigated how to reason about safety for the sampled-data system with CBFs, using a safe-triggering strategy [14] or robustifying the conventional CBF condition [15, 16, 17, 18]. On the other end of the spectrum, discrete-time CBFs were proposed in [19] for addressing safety-critical control of discrete-time dynamical systems. The discrete-time CBF condition compares the difference in the CBF values at every state transition. Inspired by this formulation, directly comparing the CBF values at sampling instants was also applied to sampled-data systems in [17, 20]. However, a direct application of a discrete-time CBF condition does not provide a theoretic safety guarantee for trajectories during the period between the two sampling instants. Moreover, as will be clarified below, if the approximate discrete-time model is not well-chosen for a continuous-time system, then a high-relative degree constraint for that system will remain high-relative degree for the discrete-time system. This phenomenon would require the discrete-time high-order CBF technique [21] to derive a valid CBF condition, further complicating the safety-critical control design.
@@
Notation: Let $\mathbb{R}$ and $\mathbb{R}_{+}$ respectively denote the reals and non-negative reals. A continuous-time signal $\bm{x}:\mathbb{R}\to\mathbb{R}^{n}$ is denoted with boldface. A continuous function $\gamma:\mathbb{R}\to\mathbb{R}$ is an extended class $\mathcal{K}$ function if it is strictly increasing and $\gamma(0)=0$.
@@ II Preliminaries and problem formulation
where $u_{k}$ is constant between consecutive sampling instants $t_{k}$ and $t_{k+1}$. We assume a uniform sampling strategy with a sampling interval $T$. As the system is piece-wise locally Lipschitz continuous, it admits a unique trajectory. Denote by $\phi(t;x,\bm{u})$ the trajectory starting at $x$ under input signal $\bm{u}$ and evaluated at time $t$. When the input signal is constant, i.e., $\bm{u}(\tau)=v$ for $\tau\in[0,t)$, we also denote it by $\phi(t;x,v)$.
@@
To accommodate the state-input dependent safety constraint (3) using a CBF, common approaches include a state projection-based approach and an integral CBF approach [12]. The former approach constructs a CBF for the projected safe state set $\mathcal{C}_{S}=\{x\in\mathbb{R}^{n}:\exists u\in U\textup{ s.t. }h(x,u)\geq 0\}$, derives relevant CBF conditions, and imposes the constraint $h(x,u)\geq 0$ as a nonlinear constraint, together with the CBF condition. When these conditions are met by a control input, the feedback renders the projected state set $\mathcal{C}_{S}$ forward invariant. The latter approach computes a virtual input for a dynamically extended system. In general, the former approach requires to construct a CBF, solve nonlinear constrained optimization, and may encounter a compatibility issue for two constraints. The integral CBF approach does not provide a clear relation between the actual input and its effects on the safe constraint function value.
In this paper, we aim to develop a new CBF methodology for the sampled-data system in (1)-(2) that handles both high-order constraints as well as state-input dependent constraints in a straightforward manner, and provides safety guarantee at all times.
@@ III Zero-order control barrier functions
A continuous safety constraint function $h$ is a zero-order control barrier function (ZOCBF) for the sampled-data system (1)-(2) with a sampling time interval $T$, if for any ${(x,u^{-})\in\mathbb{R}^{n}\times U}$, there exists an input $u\in U$ such that:
$h(\phi(T;x,u),u)-h(x,u^{-})\geq-\gamma(h(x,u^{-}))+\delta,$ (4)
@@
When interpreting the ZOCBF condition at sampling time $t_{k}$, the state-input pair $(x,u^{-})$ represents the state at $t_{k}$ and the input from an infinitesimal moment ago, i.e., the constant input $u_{k-1}$ applied during the time interval $[t_{k-1},t_{k})$. Thanks to a continuity property, $\lim_{t\to t_{k}^{-}}h(\bm{x}(t),\bm{u}(t))=h(x_{k},u_{k-1})$. This ZOCBF condition limits the change in the values of $h$ at consecutive sampling instants $t_{k}$ and $t_{k+1}$. The parameter $\delta$ serves as a robustness buffer to account for the sampling effects, as shown below. We note that the ZOCBF condition in (4) can be restricted to the set $\mathcal{C}$ when only the forward invariance of the safe set is concerned.
@@
Will the control input $u$ always appear in ZOCBF condition (4)? In the conventional continuous-time CBF literature [2], when the constraint has a high-relative degree, i.e., when $L_{g}h(x)=0$ for all $x\in\mathbb{R}^{n}$, the input will not appear in the CBF condition, rendering it non-binding. Particular techniques are proposed to address this issue, including exponential[8], high-order[9, 10] and backstepping CBFs [11]. For the proposed ZOCBF condition (4), when the safety constraint $h$ explicitly depends upon the control input, it will naturally appear. Consider the case when the constraint function $h$ does not explicitly depend on the system input, but imposes a high-relative-degree constraint. Below, we show that the input will appear in our ZOCBF condition.
@@
where $r(x)=L_{g}L_{f}^{\rho-1}y(x)\neq 0,a(x)=L_{f}^{\rho}y(x)$. More details can be found in [23, Chapter 13.2]. Now consider the ZOCBF condition (4) at a sampling time $t_{k}$. Since the dynamics of $\xi$ form a chain of integrators and $r(x)\neq 0$ (due to $h$ having relative degree $\rho$), input $u$ will appear in $\xi_{1}(t_{k+1})$, or equivalently, in $h(\bm{x}(t_{k+1}),\cdot)$. We thus conclude that $u$ will appear in the ZOCBF condition (4).
@@
Discrete-time CBFs: The simplified condition (8) seems exactly the same as the discrete-time CBF condition in [19]. Compared to a discrete-time CBF, by adding the robustness term $\delta$, the ZOCBF condition (4) guarantees system safety throughout the sampling intervals, while the discrete-time CBF condition only guarantees safety at the sampling time instants. It is important to note that for some discrete-time systems and constraints, the constraints may have a high-relative degree in the discrete-time formulation. In these cases, one must resort to applying high-order discrete-time CBFs [21], which requires the control system to advance/backtrack by multiple sampling steps before the system input appears. In strong contrast, we have shown that the system input will always appear in our proposed ZOCBF condition, regardless of the safety constraint’s relative degree. For numerical implementations, this feature implies that we need not worry about high-relative degree constraints in the ZOCBF framework if we can “accurately” predict the ZOCBF value at the next sampling instant. This difference, as well as implementation details, will be discussed in Section IV.B and demonstrated in a numerical example in Section V.A.
@@
There are two reasons why a basic safety constraint function $h(\cdot)$ may not be suitable for an online optimization-based controller (e.g., a model predictive control (MPC)[24]). First, this vanilla safety constraint may be nonlinear, and thus not suitable for constraining an online optimization problem, especially when it is used in each step of a multiple-step look-ahead planner, such as MPC. On the contrary, our ZOCBF condition can be implemented as a convex, or even linear, constraint–see the next section. Even if the ZOCBF is nonlinear, safety is guaranteed using only one-step look-ahead prediction. Second, when directly embedded into an online nonlinear MPC, the vanilla constraint starts to influence the system only when the (predicted) trajectory reaches the safety boundary, thus resulting in aggressive maneuvers with small safety margins. The influence of the safety constraint goes into effect much earlier if the ZOCBF condition is instead applied, leading to smoother and safer trajectories.
@@ IV Safe control and numerical implementations
This section demonstrates how to apply our proposed ZOCBF condition as an online safety filter. Similar to conventional CBFs, for given input signal $\bm{u}_{\text{nom}}:\mathbb{R}_{+}\to\mathbb{R}^{m}$, the filtered safe input is implemented via online optimization:
@@
This optimization yields a control input that is safe and has the least deviation from the nominal control input.
@@ IV-A Approach 1: dynamics linearization
Define the following notations. At sampling time $t_{k}$, let $x_{k}$ be the the current state, $u,u_{k-1}$ the inputs for the durations $[t_{k},t_{k+1}),[t_{k-1},t_{k})$, respectively. We propose to use a linear model to approximate the dynamics around the state $x_{k}$:
@@
where $A:=\frac{\partial f}{\partial x}(x_{k}),B:=g(x_{k}),C:=f(x_{k})-\frac{\partial f}{\partial x}(x_{k})x_{k}$. Unlike most dynamics linearization settings, here $x_{k}$ need not be an equilibrium point. Integrating the continuous-time time-invariant system (11) yields the solution
@@ IV-B Approach 2: numerical integration
We can alternatively use numerical integration methods (forward Euler, Runge-Kutta (RK), etc) to approximate the flow map. One can generally tune the desired precision of the approximate flow at the expense of computation burden. Compared to dynamic linearization, numerical integration does not use differentiation of the nonlinear vector fields, but instead uses function evaluations. This approach usually yields an explicit yet nonlinear relation between the input $u_{k}$ in the time interval $[t_{k},t_{k+1})$ and the predicted state $\hat{x}_{k+1}$ of the safety constraint at $t_{k+1}$. In the following we show that with proper integration methods, the input will appear in the ZOCBF condition (4).
We note that for input-dependent CBFs, $h(x,u)$, the input $u$ appears in (4) naturally. We restrict our discussion to state-dependent constraints $h(x,\cdot)$ here. To explore the relationship between the order of numerical integration methods and the appearance of $u$ in the ZOCBF condition, we demonstrate the results for a widely used RK integration method. As for the continuous-time systems, we consider the transformed dynamics in its normal form as in (7) to facilitate our discussion. The $p$-th order RK integration for an integration duration $T$ is given in the general form:
@@
Here $[\cdot]_{j}:(x_{1},\ldots,x_{n})\mapsto x_{j}$ is an operation that outputs the $j$th element from an vector. We apply the transformed dynamics to obtain the second equality. From (19), we know $[\varrho_{i}]_{2}=z_{2}(t_{k})+T\sum_{j=1}^{i-1}\lambda_{i,j}[\varrho_{j}]_{3}.$ Continuing recursively for $\sigma=3,\ldots,\rho-1$ yields:
@@
Thus, $u$ appears in $z_{1}(t_{k+1})$ for a constraint with relative degree $\rho$ only when we apply a $p$th-order RK integration with ${p\!\geq\!\rho}$. In this case, the control input appears explicitly in $h(\bm{x}(t_{k+1}))$ and influence its value, or equivalently, the ZOCBF condition.
@@ IV-C Approach 3: parallel simulation
For some applications, the system dynamics and the safety constraints may be too complex to analyze analytically. Instead of using an explicit dynamical equation, as in (1), we may have access to a simulator that can compute the flow map and evaluate the function $h$ in real time. The Gazebo, MuJoCo [25], Isaac Gym [26], Brax [27] simulation environments use physics engine and robot URDF description files to parallelize simulations of robotic movements. In other cases, we may have access to a physics informed neutral network (PINN) [28] that is well-trained from input-output data, which can be evaluated quickly. In these scenarios, one can sample the control set $U$, run a large quantity of parallel simulations using the sampled inputs in batch, and find an input that satisfies the ZOCBF constraint in (4) and minimizes the cost function in (10). Therefore, finite-sample simulations can provide exact verification of safety constraints, with a high confidence that depends upon the sampling density [29].
@@ IV-D Discussions
All numerical implementations will create discrepancies between the approximate and exact ZOCBF conditions. Assume that the discrepancy is bounded for by function $m$:
@@
This condition sufficiently implies the ZOCBF condition in (4), thus any feedback controller enforcing this condition guarantees system safety. We note that $m(T,u)$ is problem- and approach-specific. For example, if the constraint function $h:(x,u)\mapsto h(x,u)$ is Lipschitz continuous in $x$ with Lipschitz constant $c_{0}$ uniformly in $u$, and the fourth-order RK integration method is applied to approximate the flow, then there exists a constant $K>0$ such that $m(T,u)=Kc_{0}T^{5}$. When the sampling-time $T$ is small and fixed, $m(T,u)$ is a very small constant.
@@
Different approaches are suitable for different application scenarios. Dynamic linearization is most useful when the system dynamics is nearly linear. This approach provides a linear or convex quadratic constraint (supposing that the constraint function is concave) on the input. It also requires the least computation. Numerical integration is suitable for general nonlinear dynamics models, and we can tune the precision of the predicted state via the choice of integration methods. We have shown that, for high-relative degree constraints, high-order numerical integration methods yield a constraint in which the input $u$ appears. In practice, we can derive this constraint analytically using a symbolic toolbox, and include it into a nonlinear solver. The sampling-based parallel simulation approach can be used when the system dynamics are complex or unknown. It however requires the most computational resources.
Implementation details and comparisons of the different techniques are discussed next.

### v2

@@ I Introduction
However, there exist limitations to the conventional CBF approach. Safety constraints must be independent of the system input, and the function defining the safe set must have relative degree $1$. Otherwise, the CBF condition is either not well-defined or becomes non-binding. In order to accommodate these cases, several extensions to conventional CBFs have been proposed. Notably, exponential CBFs [8] and high-order CBFs [9, 10] are proposed for addressing the high-relative degree issue. However, checking whether the recursively defined safety function is a valid CBF in this approach is difficult in general, and the performance is highly sensitive to the chosen coefficients. Alternatively, back-stepping CBFs [11] define barrier functions based on virtual inputs and tracking errors. However, back-stepping can encounter an “explosion of terms” problem when the relative degree is large. On the other hand, when the safety constraints depend both on system states and system inputs, integral CBFs [12] and control-dependent CBFs [13] extend the system dynamics via a new virtual input, treat the original state and input as the new state, and then apply the conventional CBF approach for this extended system. For all these extensions, the relation between the input and its effects on the safety constraints is not explicit.
Digital implementation of a controller designed for continuous-time systems usually involves a zero-order hold, resulting in a sampled-data system formulation: the state evolves in continuous-time, while the control input is updated at discrete sampling instants, and held constant between those instants. Many works have investigated how to reason about safety for the sampled-data system with CBFs, using a safe-triggering strategy [14] or robustifying the conventional CBF condition [15, 16, 17, 18]. On the other end of the spectrum, discrete-time CBFs were proposed in [19] for addressing safety-critical control of discrete-time dynamical systems. The discrete-time CBF condition compares the difference in the CBF values at every state transition. Inspired by this formulation, directly comparing the CBF values at sampling instants was also applied to sampled-data systems in [17, 20]. However, a direct application of a discrete-time CBF condition does not provide a theoretic safety guarantee for trajectories during the period between the two sampling instants. Moreover, as will be clarified later in the paper, if the approximate discrete-time model is not well-chosen for a continuous-time system, then a high-relative degree constraint for that system will remain a high-relative degree for the discrete-time system. This phenomenon would require the discrete-time high-order CBF technique [21] to derive a valid CBF condition, further complicating the safety-critical control design.
@@
Notation: Let $\mathbb{R}$ and $\mathbb{R}_{+}$ denote the reals and non-negative reals. A continuous-time signal $\bm{x}:\mathbb{R}\to\mathbb{R}^{n}$ is denoted with boldface. A continuous function $\gamma:\mathbb{R}\to\mathbb{R}$ is an extended class $\mathcal{K}$ function if it is strictly increasing and $\gamma(0)=0$.
@@ II Preliminaries and problem formulation
where $u_{k}$ is constant between consecutive sampling instants $t_{k}$ and $t_{k+1}$. We assume a uniform sampling strategy with a sampling interval $T$. As the system is piece-wise locally Lipschitz continuous, it admits a unique trajectory. Denote by $\phi(t;x,\bm{u})$ the point reached by the trajectory starting at $x$ under input signal $\bm{u}$ at time $t$. When the input signal is constant, i.e., $\bm{u}(\tau)=v$ for $\tau\in[0,t)$, we also denote it by $\phi(t;x,v)$.
@@
To accommodate the state-input dependent safety constraint (3) using a CBF, common approaches include a state projection-based approach and an integral CBF approach [12]. The former approach constructs a CBF for the projected safe state set $\mathcal{C}_{S}=\{x\in\mathbb{R}^{n}:\exists u\in U\textup{ s.t. }h(x,u)\geq 0\}$, derives relevant CBF conditions, and imposes the constraint $h(x,u)\geq 0$ as a nonlinear constraint, together with the CBF condition. When these conditions are met by a control input, the feedback renders the projected state set $\mathcal{C}_{S}$ forward invariant. The latter approach computes a virtual input for a dynamically extended system. In general, the former approach involves constructing a CBF, solving a nonlinear constrained optimization problem, and potentially encountering constraint incompatibilities. The integral CBF approach does not provide a clear relation between the actual input and its effects on the safe constraint function value.
In this paper, we aim to develop a new CBF methodology for the sampled-data system in (1)-(2) that handles both high-order constraints as well as state-input dependent constraints in a straightforward manner and provides safety guarantees at all times.
@@ III Zero-order control barrier functions
A continuous safety constraint function $h$ is a zero-order control barrier function (ZOCBF) for the sampled-data system (1)-(2) with a sampling time interval $T$, if for any ${(x_{0},u_{0})\in\mathbb{R}^{n}\times U}$, there exists an input $u\in U$ such that:
$h(\phi(T;x_{0},u),u)-h(x_{0},u_{0})\geq-\gamma(h(x_{0},u_{0}))+\delta,$ (4)
@@
When interpreting the ZOCBF condition at sampling time $t_{k}$, the state-input pair $(x_{0},u_{0})$ in Definition 1 represents the state at $t_{k}$ and the input from an infinitesimal moment ago, i.e., the constant input $u_{k-1}$ applied during the time interval $[t_{k-1},t_{k})$. Thanks to a continuity property, $\lim_{t\to t_{k}^{-}}h(\bm{x}(t),\bm{u}(t))=h(x_{k},u_{k-1})$. This ZOCBF condition limits the change in the values of $h$ at consecutive sampling instants $t_{k}$ and $t_{k+1}$. The parameter $\delta$ serves as a robustness buffer to account for the sampling effects, as shown below. We note that the ZOCBF condition in (4) can be restricted to the set $\mathcal{C}$ when only the forward invariance of the safe set is concerned.
@@
Will the control input $u$ always appear in ZOCBF condition (4)? In the conventional continuous-time CBF literature [2], when the constraint has a high-relative degree, i.e., when $L_{g}h(x)=0$ for all $x\in\mathbb{R}^{n}$, the input will not appear in the CBF condition, rendering it non-binding. Particular techniques are proposed to address this issue, including exponential[8], high-order[9, 10] and backstepping CBFs [11]. For the proposed ZOCBF condition (4), when the safety constraint $h$ explicitly depends upon the control input, it will naturally appear. Consider the case when the constraint function $h$ does not explicitly depend on the system input but imposes a high-relative-degree constraint. Below, we show that the input appears in our ZOCBF condition.
@@
where $r(x):=L_{g}L_{f}^{\rho-1}y(x)\neq 0,a(x):=L_{f}^{\rho}y(x)$. More details can be found in [23, Chapter 13.2]. Now consider the ZOCBF condition (4) at a sampling time $t_{k}$. Since the dynamics of $\xi$ form a chain of integrators and $r(x)\neq 0$ (due to $h$ having relative degree $\rho$), input $u$ will appear in $\xi_{1}(t_{k+1})$, or equivalently, in $h(\bm{x}(t_{k+1}),\cdot)$. We thus conclude that $u$ will appear in the ZOCBF condition (4).
@@
Discrete-time CBFs: The simplified condition (8) seems exactly the same as the discrete-time CBF condition in [19]. Compared to a discrete-time CBF, by adding the robustness term $\delta$, the ZOCBF condition (4) guarantees system safety throughout the sampling intervals, while the discrete-time CBF condition only guarantees safety at the sampling time instants. From Lemma 1, $\delta$ is small if the sampling rate is fast. It is important to note that, for some discrete-time systems and constraints, the constraints may have a high-relative degree in the discrete-time formulation. In these cases, one must resort to applying high-order discrete-time CBFs [21], which requires the control system to advance/backtrack by multiple sampling steps before the system input appears. In strong contrast, we have shown that the system input will always appear in our proposed ZOCBF condition, regardless of the safety constraint’s relative degree. For numerical implementations, this feature implies that we need not worry about high-relative degree constraints in the ZOCBF framework if we can “accurately” predict the ZOCBF value at the next sampling instant. This difference, as well as implementation details, will be discussed in Section IV.B and demonstrated in a numerical example in Section V.A.
@@
There are two reasons why a basic safety constraint function $h(\cdot)$ may not be suitable to be used as it is for an online optimization-based controller (e.g., a model predictive control (MPC) [24]). First, this vanilla safety constraint may be nonlinear, and thus not suitable for constraining an online optimization problem, especially when it is used in each step of a multiple-step look-ahead planner, such as MPC. On the contrary, our ZOCBF condition can be implemented as a convex, or even linear, constraint–see the next section. Even if the ZOCBF is nonlinear, safety is guaranteed using only one-step look-ahead prediction. Second, when directly embedded into an online nonlinear MPC, the vanilla constraint starts to influence the system only when the (predicted) trajectory reaches the safety boundary, thus resulting in aggressive maneuvers with small safety margins. The influence of the safety constraint goes into effect much earlier if the ZOCBF condition is instead applied, leading to smoother and safer trajectories.
@@ IV Safe control and numerical implementations
This section demonstrates how to apply our proposed ZOCBF condition as an online safety filter. Similar to conventional CBFs, for a given nominal input signal $\bm{u}_{\text{nom}}:\mathbb{R}_{+}\to\mathbb{R}^{m}$, the filtered safe input is implemented via online optimization:
@@
This optimization yields a control input that is safe and has the least deviation from the nominal control input. The nominal input signal may come from different schemes, for example, a state feedback controller or a human operator.
@@ IV-A Approach 1: dynamics linearization
Define the following notations. At sampling time $t_{k}$, let $x_{k}$ be the current state, $u,u_{k-1}$ the inputs for the durations $[t_{k},t_{k+1}),[t_{k-1},t_{k})$, respectively. We propose to use a linear model to approximate the dynamics around the state $x_{k}$:
@@ IV-B Approach 2: numerical integration
We can alternatively use numerical integration methods (forward Euler, Runge-Kutta (RK), etc) to approximate the flow map. One can generally tune the desired precision of the approximate flow at the expense of computation burden. Compared to dynamic linearization, numerical integration does not use differentiation of the nonlinear vector fields but instead uses function evaluations. This approach usually yields an explicit yet nonlinear relation between the input $u_{k}$ in the time interval $[t_{k},t_{k+1})$ and the predicted state $\hat{x}_{k+1}$ of the safety constraint at $t_{k+1}$. In the following, we show that by substituting the predicted state $\hat{x}_{k+1}$ in the ZOCBF condition (4), the input will appear. For notational simplicity, we ignore the small discrepancy between the predicted and the exact states in this subsection, which will be accounted for in a following subsection.
For state-input dependent CBFs, $h(x,u)$, the input $u$ appears in (4) naturally. We restrict our discussion to state-dependent constraints $h(x,\cdot)$ here. To explore the relationship between the order of numerical integration methods and the appearance of $u$ in the ZOCBF condition, we demonstrate the results for a widely used RK integration method. As for the continuous-time systems, we consider the transformed dynamics in its normal form as in (7) to facilitate our discussion. The $p$-th order RK integration [20] for an integration duration $T$ is given in the general form:
@@
Here $[\cdot]_{j}:(x_{1},\ldots,x_{n})\mapsto x_{j}$ is an operation that outputs the $j$th element from a vector. We apply the transformed dynamics to obtain the second equality. From (18), we know $[\varrho_{i}]_{2}=z_{2}(t_{k})+T\sum_{j=1}^{i-1}\lambda_{i,j}[\varrho_{j}]_{3}.$ Continuing recursively for $\sigma=3,\ldots,\rho-1$ yields:
@@
Thus, $u$ appears in $z_{1}(t_{k+1})$ for a constraint with relative degree $\rho$ only when we apply a $p$th-order RK integration with ${p\!\geq\!\rho}$. In this case, the control input appears explicitly in $h(\bm{x}(t_{k+1}))$ and influences its value, or equivalently, the ZOCBF condition.
@@ IV-C Approach 3: parallel simulation
For some applications, the system dynamics and the safety constraints may be too complex to analyze analytically. Instead of using an explicit dynamical equation, as in (1), we may have access to a simulator that can compute the flow map and evaluate the function $h$ in real time. Gazebo, MuJoCo [25], and Isaac Gym [26], and Brax [27] simulation environments use physics engine and robot URDF description files to parallelize simulations of robotic movements. In other cases, we may have access to a physics-informed neural network (PINN) [28] that is well-trained from input-output data, which can be evaluated quickly. In these scenarios, one can sample the control set $U$, run a large number of parallel simulations using the sampled inputs in batch, and find an input that satisfies the ZOCBF constraint in (4) and minimizes the cost function in (10). Therefore, finite-sample simulations can provide exact verification of safety constraints, with high confidence, depending on the sampling density [29].
@@ IV-D Discussions
All numerical implementations will create discrepancies between the approximate and the exact ZOCBF conditions. Assume that the discrepancy is bounded by function $m$:
@@
This condition sufficiently implies the ZOCBF condition in (4), thus any feedback controller enforcing this condition guarantees system safety. We note that $m(T,u)$ is problem- and approach-specific. For example, if the constraint function $h:(x,u)\mapsto h(x,u)$ is Lipschitz continuous in $x$ with Lipschitz constant $c_{0}$ uniformly in $u$, and the fourth-order RK integration method is applied to approximate the flow, then there exists a constant $K>0$ such that $m(T,u)=Kc_{0}T^{5}$. When the sampling time $T$ is small and fixed, $m(T,u)$ is a fairly small constant.
@@
Different approaches are suitable for different application scenarios. Dynamic linearization is most useful when the system dynamics is nearly linear. This approach provides a linear or convex quadratic constraint (supposing that the constraint function is concave) on the input. It also requires the least computation. Numerical integration is suitable for general nonlinear dynamics models, and we can tune the precision of the predicted state via the choice of integration methods. We have shown that, for high-relative degree constraints, high-order numerical integration methods yield a constraint in which the input $u$ appears. In practice, we can derive this constraint analytically using a symbolic toolbox, and include it into a nonlinear solver. The sampling-based parallel simulation approach can be used when the system dynamics are complex or unknown. It however requires the most computational resources. Implementation details and comparisons of different techniques are discussed next.